Circular No. 6356

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744–28 D. M. Cole, D. E. Vanden Berk, R. C. Nichol, J. M. Quashnock, S. A.

Severson, M. C. Miller, D. Q. Lamb, University of Chicago, report: “We have determined the HST Guide Star reference frame positions of seven bright stars encircling the position of the variable radio source (IAUC 6307, 6323) lying within the XTE error box (IAUC 6291) for GRO J1744–28, using the STScI Digitized Sky Survey scans and the IRAF/STSDAS GASP package. Using these stars (finding charts of which are available from us upon request), we have calculated the rms positions and errors of two stars in the optical and five stars in the infrared near the position of the variable radio source on images obtained with the help of R. Rutledge (the coordinates are end figures for α = 17^{h}44^{m} and δ = −28^{o}45′; equinox 2000.0): star ‘a’, 36^{s}.844 ± 0^{s}.004, 38′′.16 ± 0′′.08; star ‘b’, 36^{s}.655 ± 0^{s}.010, 36′′.93 ± 0′′.10; star ‘A’, 37^{s}.035 ± 0^{s}.004, 37′′.59 ± 0′′.08; star ‘B’, 37^{s}.068 ± 0^{s}.004, 35′′.91 ± 0′′.07; star ‘C’, 36^{s}.844 ± 0^{s}.005, 38′′.28 ± 0′′.08; star ‘D’, 36^{s}.998 ± 0^{s}.004, 34′′.26 ± 0′′.07; star ‘E’, 36^{s}.640 ± 0^{s}.005, 37′′.98 ± 0′′.08. The astrometry was performed on images obtained from IAUC 6310, 6314, and 6326; and private communication from Augusteijn and Van de Steene; Fletcher, Cooray, and Burke; Blanco, Lidman, and Glazebrook; and Keane. The detections of star ‘a’ in the UKST B_{j} (IAUC 6314) and the ESO NTT V (IAUC 6326) images are marginal, so we did not use those images in our calculation of the position of this star. Star ‘b’ is the second star in the I and R reported by Augusteijn et al. (IAUC 6326), and star ‘E’ is the star immediately west of ‘C’ in the near-infrared images. We note that there are no other objects visible in the vicinity of the variable radio source that are brighter than approximately 22 in R (IAUC 6326) and 22 in J (Keane, private comm.). Although our relative astrometry is excellent, our absolute astrometry is limited by systematic offsets between the HST Guide Star and radio coordinate systems. In order to estimate the offset for this field, we have compared the 17 PPM stars within 0.7 degrees of the radio source with their positions in the HST Guide Star Catalogue. We find α(PPM– HST) = −0′′.333 ± 0′′.295, δ(PPM–HST) = −0′′.024 ± 0′′.285. Star ‘a’ and ‘C’ are almost certainly the same star, but our position for this star is 0′′.9 away from the position of the variable radio source. It is not clear that the systematic offset between the HST Guide Star and radio coordinate systems is sufficient to reconcile the two positions. All other stars visible in the optical or infrared lie more than 2′′ from the radio source.”

1996 March 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green